# Towards Automated Identification of Violation Symptoms of Architecture Erosion

### v3

@@ 3 Study Design
Figure 2 provides an overview of the whole process. As illustrated in this figure (see the top box, Dataset Creation), we collected data and established a dataset on architecture violations in our previous work [11]. During Phase 1, we explored the possibility of using this dataset to train ML/DL-based classifiers to automatically identify violation symptoms in code review comments. During Phase 2, we sent the survey questionnaire to 200 involved participants and received 24 responses that were used to validate the usefulness of automated identification of violation symptoms in practice.
Figure 2: An overview of the research process
3.1 Research Questions
In this work, we aim to explore the feasibility of automatically identifying violation symptoms of architecture erosion from code review comments. Specifically, we formulated the goal of this study by following the Goal-Question-Metric method [33]: analyze code review comments for the purpose of exploring the feasibility of automatically identifying violation symptoms of architecture erosion with respect to the performance on architecture erosion identification from the point of view of developers in the context of open source software development. To achieve our goal, we formulate three Research Questions (RQs):
@@ 3.1 Research Questions
Rationale: This RQ focuses on training feasible classifiers to automatically identify violation symptoms from textual artifacts. As mentioned in Section 1, the percentage of violation symptoms (belonging to architecture-level issues) is relatively lower than code-level issues. We thus expect that the description of violation symptoms is scattered in textual artifacts during development. RQ1.1 focuses on finding out the classifier with the best performance to distinguish violation symptoms from natural language by comparing several popular ML and DL algorithms. RQ1.2 aims at employing several well-known word embedding models to generate word vectors as features of labeled data, in order to choose the word embedding models that can help to improve the accuracy of classifiers. The word embedding models are a common ground for both ML and DL-based classifiers in our study. Finally, with RQ1.3, we seek to compare the differences in dimension size among different word embedding models, as it is one of the most important hyperparameters affecting the model performance[34]. Specifically, RQ1.3 endeavors to shed light on the extent to which the selection of dimensionality impacts the performance of the classifiers in automatically identifying violation symptoms.
@@ 3.2 Data Collection
We relied on the dataset that we mined and manually identified in our previous study [11]. As shown in Figure 2 (see Dataset Creation), we followed five steps to establish a specific dataset on architecture violations mined from code review comments. More specifically, we mined code review comments of four OSS projects in the OpenStack (i.e., Nova and Neutron) and Qt (i.e., Qt Base and Qt Creator) communities between 2014 and 2020. We mined the data through the REST API [^4] https://gerrit-review.googlesource.com/Documentation/rest-api.html supported by Gerrit. We first conducted experiment-based keyword refinement to improve our keyword search, as our previous experience indicated that a keyword-based approach could only capture very small proportions of code review comments containing violation symptoms [11]. Subsequently, we proceeded with manual identification, and pilot and formal data labeling of review comments related to architecture violations. In total, we manually identified and labeled 606 code review comments pertaining to architecture violations. Code review comments were created for patches, and comments were made by code reviewers and developers. In terms of the data labeling process, four researchers first conducted a pilot data labeling by randomly selecting 50 review comments, to reach a consensus (the inter-rater agreement between the researchers measured in the Cohen’s Kappa coefficient value [35] was 0.857) and ensure that we have the same understanding of violation symptoms. Any disagreements were discussed between the four researchers to reach a consensus. After that, the four researchers started the formal data labeling by dividing the retrieved 21,583 code review comments into four parts (each researcher manually labeled around 5,400 comments). After the formal data labeling, the first author checked the data labeling results from the other three researchers to make sure that there were no false positives. To mitigate potential bias, the four researchers discussed all the conflicts in the labeling results until we reached an agreement. To summarize, all data labeling results were checked by at least two researchers and conflicts were discussed among four researchers. Subsequently, all the collected code review comments regarding architecture violations (i.e., 606 code review comments) from the four projects were combined into an integrated dataset used in this work. The dataset has been made available in a replication package [36].
3.3 Phase 1 - Automatic Classification of Violation Symptoms
@@
In this section, we describe the experimental setup that we followed to evaluate the performance of the trained classifiers (see Figure 3). The steps in this setup correspond to the steps of Phase 1 in Figure 2. First, we pre-processed the collected code review comments to generate structured word sets, and then utilized word embedding techniques to encode the words and generate vectors for presenting the words in review comments. Subsequently, the generated vectors acted as the input to train the classifiers based on ML/DL models, in order to learn to classify review comments as violations and non-violations. The experimental environment is a computer equipped with Intel(R) Core(TM) i7-10510U CPU and 16GB RAM, running Windows 11 (64-bit). The next subsections elaborate on these steps for both the ML and DL models. More details of the experiments (such as hyperparameters and experimental environment) can be found in the replication package [36].
@@
Figure 3: The framework of the experimental setup for classifying violation symptoms of architecture erosion from code review comments
3.3.1 Machine Learning Models
@@
Step 3. Stop words Removal: Stop words are frequently observed in text, for the purpose of distinguishing between texts, such as “the”, “are”, and “is”, but they are typically devoid of valuable semantic content. Thus, we removed stop words in our dataset by using the Natural Language Toolkit (NLTK) [37].
@@
Step 5. Stemming: Stemming can be used to convert the words to their base forms. For instance, “architecture” and “architectural” have the same stem (i.e., token) “architectu”. We conducted a stemming process (using Snowball Stemmer from the NLTK toolkit [37]) to obtain the stem of each token, since it is a more comprehensive and aggressive version of the Porter Stemmer [38].
@@
  • word2vec: word2vec was developed in 2013 [39] and it takes words as input and provides a word embedding matrix that contains high-dimensional unique vectors to represent each distinct word in given training corpora. word2vec can generate word embeddings through skip-gram and Continuous Bag-of-Words (CBOW) models. In this study, we adopted the pre-trained word2vec model proposed by Efstathiou et al. [40]. This model was trained based on the corpus comprising over 15 GB of textual data collected from Stack Overflow posts, which contains a plethora of textual information in the software engineering domain.
  • fastText: fastText is an open-source and lightweight library proposed by Meta AI Research lab in 2017 [41] and it is capable of learning text representations and training text classifiers. Based on the fastText model, Meta published pre-trained word vectors with 300 dimensions for 157 languages constructed on Common Crawl and Wikipedia [^5] https://fasttext.cc/docs/en/crawl-vectors.html. In this study, we employed the pre-trained fastText model, which can be transformed into 100 and 200-dimensional models, in order to control variables and avoid the impact of dimensionality on classifiers’ performance.
  • GloVe: Global Vectors for Word Representation (GloVe) is another common word embedding technique [42]. A GloVe model is inspired by aggregated word co-occurrence statistics that may encode global information for words. Compared with word2vec, GloVe avoids the weakness regarding using local word co-occurrence information. In this study, we adopted the GloVe pre-trained word vectors on Twitter data (two billion tweets) with 200 dimensions.
Model Training. Given the binary classification task in this study, we selected five common ML algorithms that are widely used in classification tasks [43], including Support Vector Machine (SVM), Logistic Regression (LR), Decision Tree (DT), Bernoulli Naive Bayes (NB), and k-Nearest Neighbor (kNN). These classification algorithms are also commonly applied in many areas of software engineering studies [44].
For the ML classifiers, we performed a grid search to tune the hyperparameters of each model and picked the best-performing ML models with specific hyperparameters on the validation set. Grid search, an exhaustive search technique, explores every possible combination within a predefined search space to identify the optimal configuration. More specifically, this method is widely employed in the realms of machine learning and optimization and it involves a systematic evaluation that examines all of the combinations of a set of candidate settings to find the best hyperparameter combination [34].
3.3.2 Deep Learning Models
@@
Model Training. In recent years, DL-based models are increasingly used in various text classification tasks and Convolutional Neural Network (CNN) has become one of the most popular model architectures for text classification [45]. We selected the TextCNN model to train the DL-based classifiers, since TextCNN, as an adapted CNN-based model, is reported to improve the state of the art on text classification [45]. TextCNN is a CNN-based architecture designed for processing textual data, and was introduced by Yoon Kim in 2014 [46]. It can be used for a variety of Natural Language Processing (NLP) tasks, such as text classification and sentiment analysis. Compared with traditional ML models regarding text processing, the key advantage of TextCNN is its ability to automatically learn features from raw textual data, without relying on manual feature engineering.
@@
TextCNN is a convolutional neural network model that is commonly used for text classification tasks. This model takes a sentence as input and applies multiple convolutional filters to extract important features from the text. A TextCNN classifier based on the vocabulary set of a dataset uses a one-hot encoding scheme to represent words and generate the word embedding, while a TextCNN classifier can also utilize pre-trained word embedding models to capture semantic information of words. Additionally, the TextCNN model requires a proper setting to construct neural network models, such as hyperparameters and word embedding parameters. In this work, we set the embedding dimension to 200, the learning rate to 0.001, and the dropout to 0.25 to handle the overfitting, while the filter region size is (3, 4, 5). Each DL model’s batch size is set to 16, and the network is trained for 100 epochs with the early stopping strategy with the purpose of avoiding overfitting. We configured the patience parameter to 8, indicating that the training process would terminate after 8 consecutive epochs with no observed improvement in model performance. These approaches are common practices in machine learning, aimed at preventing overfitting and optimizing the model’s convergence [47].
3.3.3 Performance Evaluation
@@
$\rm Precision=\frac{TP}{TP+FP}$ (1)
$\rm Recall=\frac{TP}{TP+FN}$ (2)
$\rm F_{1}-score=2\times\frac{Precision\times Recall}{Presion+Recall}$ (3)
$\rm Accuracy=\frac{TP+TN}{TP+TN+FP+FN}$ (4)
@@
To avoid the class imbalance problem [48], a data distribution problem that significantly impacts classifiers’ performance, we constructed our dataset with a balanced class distribution (i.e., the same number of review comments labeled as “violations” and “non-violations”), which avoids the poor generalization ability of classifiers. With respect to the model training and validation strategy, we employed 10-fold cross-validation when we conducted a performance evaluation of the classifiers. We divided the collected dataset into three parts, i.e., we randomly set aside 80% of the data for the training set (60%) and validation set (20%), and 20% for the test set. Besides, we set a fixed seed for reproducibility during experiments.
3.4 Ensemble Classifier
To identify architecture violations from textual artifacts (i.e., code review comments), we trained several ML and DL-based classifiers to predict the labels of new review comments. However, the trained classifiers have varying performances in predicting the labels of target artifacts. In RQ2, we asked whether composing the individual classifiers into an ensemble classifier can help improve the overall performance. This composition can be done by utilizing the ensemble learning technique, i,e., majority voting strategy [49]. Specifically, a combined classifier of several individual classifiers is generated with the purpose of building an improved classifier, which predicts an output (class) based on the highest probability of chosen class as the output [49]. This aggregating criterion is like an election process and can produce combined voting results from each classifier’s output.
@@ 3.4 Ensemble Classifier
3.5 Phase 2 - Validation: Survey and Interview
@@
To answer RQ3, we designed and executed a survey and semi-structured interviews by following the guidelines proposed by Kitchenham and Pfleeger’s personal opinion surveys [50]; in terms of the survey, we used an anonymous, web-based questionnaire to increase the response rate.
@@
(1) Survey. Surveys can be conducted in different ways, such as online questionnaires, phone surveys [51]. Our survey is an online questionnaire of the cross-sectional type, which allows us to collect data and compare differences between participants at a single point in time. We adopted self-administered questionnaires to collect responses, since (a) respondents can answer the survey at their convenience (reducing our time investment and increasing the response rate); (b) questionnaires can reduce research bias since the researcher does not have any contact with the respondents when they fill out the survey; and (c) respondents enjoy better privacy than interviews.
@@
We invited participants through customized emails (see the replication package [36]) attaching the links to their code review comments that contain violation symptoms and are identified by our approach, as well as the survey form (also see the replication package [36]). Our survey was constructed based on four parts: Problem Statement, Our approach, Survey Goal, and Survey Questions. Problem Statement introduces the research background regarding what is architecture erosion and its violation symptoms, as well as why it is important to investigate violation symptoms. Our approach describes the approach that we developed to automatically identify violation symptoms with high performance from a large number of code reviews as well as two examples of violation symptoms of architecture erosion detected by our approach from code review comments. In Survey Goal, we explain the purpose of this survey. In Survey Questions, we ask five general questions and present six statements concerning the usefulness of our approach. Note that, before the formal survey, we did a pilot survey (receiving 6 responses) with the purpose of revising and improving our survey questions and examples.
@@
The five general survey questions include demographic information (e.g., their countries and work experience). The six statements are assessed using a five-point Likert scale (Strong Agree, Agree, Neutral, Disagree, and Strongly Disagree) with one option (“I don’t know”). This scale enables participants to rate the extent that they agree or disagree with the statements. Finally, at the end of the survey, we ask an open-ended question to let the respondents freely provide comments and suggestions about our survey. To formulate the statements, we took inspiration from the survey questions of two survey studies [52, 53] that evaluate the usefulness of automatic classifiers based on ML and DL techniques in the software engineering community. We chose to use declarative statements in our validation survey since declarative statements are deemed appropriate for exploratory studies to convey simple and clear judgements; in contrast, descriptive statements are more suitable in explanatory studies to provide the details of the judgements in various aspects. The statements are as follows:
@@
Data Analysis. For the survey questions, we applied descriptive statistics to analyze the responses to the demographic and Likert scale questions. For the open-ended interview questions, interviews were recorded and transcribed; subsequently, we used the open coding and constant comparison method [54] to analyze and categorize the qualitative responses.
4 Results
4.1 RQ1: Identifying Violation Symptoms

### v4

@@ 3. Study Design
Figure 2 provides an overview of the whole process. As illustrated in this figure (see the top box, Dataset Creation), we collected data and established a dataset on architecture violations in our previous work (Li2023wvs, ). During Phase 1, we explored the possibility of using this dataset to train ML/DL-based classifiers to automatically identify violation symptoms in code review comments. During Phase 2, we sent the survey questionnaire to 200 involved participants and received 24 responses that were used to validate the usefulness of automated identification of violation symptoms in practice.
Figure 2. An overview of the research process
3.1. Research Questions
In this work, we aim to explore the feasibility of automatically identifying violation symptoms of architecture erosion from code review comments. Specifically, we formulated the goal of this study by following the Goal-Question-Metric method (Basili1994gmq, ): analyze code review comments for the purpose of exploring the feasibility of automatically identifying violation symptoms of architecture erosion with respect to the performance on architecture erosion identification from the point of view of developers in the context of open source software development. To achieve our goal, we formulate three Research Questions (RQs):
@@ 3.1. Research Questions
Rationale: This RQ focuses on training feasible classifiers to automatically identify violation symptoms from textual artifacts. As mentioned in Section 1, the percentage of violation symptoms (belonging to architecture-level issues) is relatively lower than code-level issues. We thus expect that the description of violation symptoms is scattered in textual artifacts during development. RQ1.1 focuses on finding out the classifier with the best performance to distinguish violation symptoms from natural language by comparing several popular ML and DL algorithms. RQ1.2 aims at employing several well-known word embedding models to generate word vectors as features of labeled data, in order to choose the word embedding models that can help to improve the accuracy of classifiers. The word embedding models are a common ground for both ML and DL-based classifiers in our study. Finally, with RQ1.3, we seek to compare the differences in dimension size among different word embedding models, as it is one of the most important hyperparameters affecting the model performance(hutter2019aml, ). Specifically, RQ1.3 endeavors to shed light on the extent to which the selection of dimensionality impacts the performance of the classifiers in automatically identifying violation symptoms.
@@ 3.2. Data Collection
We relied on the dataset that we mined and manually identified in our previous study (Li2023wvs, ). As shown in Figure 2 (see Dataset Creation), we followed five steps to establish a specific dataset on architecture violations mined from code review comments. More specifically, we mined code review comments of four OSS projects in the OpenStack (i.e., Nova and Neutron) and Qt (i.e., Qt Base and Qt Creator) communities between 2014 and 2020. We mined the data through the REST API [^4] https://gerrit-review.googlesource.com/Documentation/rest-api.html supported by Gerrit. We first conducted experiment-based keyword refinement to improve our keyword search, as our previous experience indicated that a keyword-based approach could only capture very small proportions of code review comments containing violation symptoms (Li2023wvs, ). Subsequently, we proceeded with manual identification, and pilot and formal data labeling of review comments related to architecture violations. In total, we manually identified and labeled 606 code review comments pertaining to architecture violations. Code review comments were created for patches, and comments were made by code reviewers and developers. In terms of the data labeling process, four researchers first conducted a pilot data labeling by randomly selecting 50 review comments, to reach a consensus (the inter-rater agreement between the researchers measured in the Cohen’s Kappa coefficient value (Cohen1960cans, ) was 0.857) and ensure that we have the same understanding of violation symptoms. Any disagreements were discussed between the four researchers to reach a consensus. After that, the four researchers started the formal data labeling by dividing the retrieved 21,583 code review comments into four parts (each researcher manually labeled around 5,400 comments). After the formal data labeling, the first author checked the data labeling results from the other three researchers to make sure that there were no false positives. To mitigate potential bias, the four researchers discussed all the conflicts in the labeling results until we reached an agreement. To summarize, all data labeling results were checked by at least two researchers and conflicts were discussed among four researchers. Subsequently, all the collected code review comments regarding architecture violations (i.e., 606 code review comments) from the four projects were combined into an integrated dataset used in this work. The dataset has been made available in a replication package (onlinepackage_TSE, ).
3.3. Phase 1 - Automatic Classification of Violation Symptoms
@@
In this section, we describe the experimental setup that we followed to evaluate the performance of the trained classifiers (see Figure 3). The steps in this setup correspond to the steps of Phase 1 in Figure 2. First, we pre-processed the collected code review comments to generate structured word sets, and then utilized word embedding techniques to encode the words and generate vectors for presenting the words in review comments. Subsequently, the generated vectors acted as the input to train the classifiers based on ML/DL models, in order to learn to classify review comments as violations and non-violations. The experimental environment is a computer equipped with Intel(R) Core(TM) i7-10510U CPU and 16GB RAM, running Windows 11 (64-bit). The next subsections elaborate on these steps for both the ML and DL models. More details of the experiments (such as hyperparameters and experimental environment) can be found in the replication package (onlinepackage_TSE, ).
@@
Figure 3. The framework of the experimental setup for classifying violation symptoms of architecture erosion from code review comments
3.3.1. Machine Learning Models
@@
Step 3. Stop words Removal: Stop words are frequently observed in text, for the purpose of distinguishing between texts, such as “the”, “are”, and “is”, but they are typically devoid of valuable semantic content. Thus, we removed stop words in our dataset by using the Natural Language Toolkit (NLTK) (Bird2010nlp, ).
@@
Step 5. Stemming: Stemming can be used to convert the words to their base forms. For instance, “architecture” and “architectural” have the same stem (i.e., token) “architectu”. We conducted a stemming process (using Snowball Stemmer from the NLTK toolkit (Bird2010nlp, )) to obtain the stem of each token, since it is a more comprehensive and aggressive version of the Porter Stemmer (SnowballStemmer, ).
@@
  • word2vec: word2vec was developed in 2013 (Mikolov2013word2vec, ) and it takes words as input and provides a word embedding matrix that contains high-dimensional unique vectors to represent each distinct word in given training corpora. word2vec can generate word embeddings through skip-gram and Continuous Bag-of-Words (CBOW) models. In this study, we adopted the pre-trained word2vec model proposed by Efstathiou et al. (Efstathiou2018weso, ). This model was trained based on the corpus comprising over 15 GB of textual data collected from Stack Overflow posts, which contains a plethora of textual information in the software engineering domain.
  • fastText: fastText is an open-source and lightweight library proposed by Meta AI Research lab in 2017 (Bojanowski2017fastText, ) and it is capable of learning text representations and training text classifiers. Based on the fastText model, Meta published pre-trained word vectors with 300 dimensions for 157 languages constructed on Common Crawl and Wikipedia [^5] https://fasttext.cc/docs/en/crawl-vectors.html. In this study, we employed the pre-trained fastText model, which can be transformed into 100 and 200-dimensional models, in order to control variables and avoid the impact of dimensionality on classifiers’ performance.
  • GloVe: Global Vectors for Word Representation (GloVe) is another common word embedding technique (Pennington2014GloVe, ). A GloVe model is inspired by aggregated word co-occurrence statistics that may encode global information for words. Compared with word2vec, GloVe avoids the weakness regarding using local word co-occurrence information. In this study, we adopted the GloVe pre-trained word vectors on Twitter data (two billion tweets) with 200 dimensions.
Model Training. Given the binary classification task in this study, we selected five common ML algorithms that are widely used in classification tasks (Han2022dm, ), including Support Vector Machine (SVM), Logistic Regression (LR), Decision Tree (DT), Bernoulli Naive Bayes (NB), and k-Nearest Neighbor (kNN). These classification algorithms are also commonly applied in many areas of software engineering studies (Yang2022pmse, ).
For the ML classifiers, we performed a grid search to tune the hyperparameters of each model and picked the best-performing ML models with specific hyperparameters on the validation set. Grid search, an exhaustive search technique, explores every possible combination within a predefined search space to identify the optimal configuration. More specifically, this method is widely employed in the realms of machine learning and optimization and it involves a systematic evaluation that examines all of the combinations of a set of candidate settings to find the best hyperparameter combination (hutter2019aml, ).
3.3.2. Deep Learning Models
@@
Model Training. In recent years, DL-based models are increasingly used in various text classification tasks and Convolutional Neural Network (CNN) has become one of the most popular model architectures for text classification (Minaee2021dltc, ). We selected the TextCNN model to train the DL-based classifiers, since TextCNN, as an adapted CNN-based model, is reported to improve the state of the art on text classification (Minaee2021dltc, ). TextCNN is a CNN-based architecture designed for processing textual data, and was introduced by Yoon Kim in 2014 (Kim2014TextCNN, ). It can be used for a variety of Natural Language Processing (NLP) tasks, such as text classification and sentiment analysis. Compared with traditional ML models regarding text processing, the key advantage of TextCNN is its ability to automatically learn features from raw textual data, without relying on manual feature engineering.
@@
TextCNN is a convolutional neural network model that is commonly used for text classification tasks. This model takes a sentence as input and applies multiple convolutional filters to extract important features from the text. A TextCNN classifier based on the vocabulary set of a dataset uses a one-hot encoding scheme to represent words and generate the word embedding, while a TextCNN classifier can also utilize pre-trained word embedding models to capture semantic information of words. Additionally, the TextCNN model requires a proper setting to construct neural network models, such as hyperparameters and word embedding parameters. In this work, we set the embedding dimension to 200, the learning rate to 0.001, and the dropout to 0.25 to handle the overfitting, while the filter region size is (3, 4, 5). Each DL model’s batch size is set to 16, and the network is trained for 100 epochs with the early stopping strategy with the purpose of avoiding overfitting. We configured the patience parameter to 8, indicating that the training process would terminate after 8 consecutive epochs with no observed improvement in model performance. These approaches are common practices in machine learning, aimed at preventing overfitting and optimizing the model’s convergence (Watson2022SLRDL, ).
3.3.3. Performance Evaluation
@@
(1) $\rm Precision=\frac{TP}{TP+FP}$
(2) $\rm Recall=\frac{TP}{TP+FN}$
(3) $\rm F_{1}-score=2\times\frac{Precision\times Recall}{Presion+Recall}$
(4) $\rm Accuracy=\frac{TP+TN}{TP+TN+FP+FN}$
@@
To avoid the class imbalance problem (he2009imbalance, ), a data distribution problem that significantly impacts classifiers’ performance, we constructed our dataset with a balanced class distribution (i.e., the same number of review comments labeled as “violations” and “non-violations”), which avoids the poor generalization ability of classifiers. With respect to the model training and validation strategy, we employed 10-fold cross-validation when we conducted a performance evaluation of the classifiers. We divided the collected dataset into three parts, i.e., we randomly set aside 80% of the data for the training set (60%) and validation set (20%), and 20% for the test set. Besides, we set a fixed seed for reproducibility during experiments.
3.4. Ensemble Classifier
To identify architecture violations from textual artifacts (i.e., code review comments), we trained several ML and DL-based classifiers to predict the labels of new review comments. However, the trained classifiers have varying performances in predicting the labels of target artifacts. In RQ2, we asked whether composing the individual classifiers into an ensemble classifier can help improve the overall performance. This composition can be done by utilizing the ensemble learning technique, i,e., majority voting strategy (Dietterich2000emml, ). Specifically, a combined classifier of several individual classifiers is generated with the purpose of building an improved classifier, which predicts an output (class) based on the highest probability of chosen class as the output (Dietterich2000emml, ). This aggregating criterion is like an election process and can produce combined voting results from each classifier’s output.
@@ 3.4. Ensemble Classifier
3.5. Phase 2 - Validation: Survey and Interview
@@
To answer RQ3, we designed and executed a survey and semi-structured interviews by following the guidelines proposed by Kitchenham and Pfleeger’s personal opinion surveys (gaese2008CH3, ); in terms of the survey, we used an anonymous, web-based questionnaire to increase the response rate.
@@
(1) Survey. Surveys can be conducted in different ways, such as online questionnaires, phone surveys (Lethbridge2005sse, ). Our survey is an online questionnaire of the cross-sectional type, which allows us to collect data and compare differences between participants at a single point in time. We adopted self-administered questionnaires to collect responses, since (a) respondents can answer the survey at their convenience (reducing our time investment and increasing the response rate); (b) questionnaires can reduce research bias since the researcher does not have any contact with the respondents when they fill out the survey; and (c) respondents enjoy better privacy than interviews.
@@
We invited participants through customized emails (see the replication package (onlinepackage_TSE, )) attaching the links to their code review comments that contain violation symptoms and are identified by our approach (i.e., the best-performing classifier), as well as the survey form (also see the replication package (onlinepackage_TSE, )). Our survey was constructed based on four parts: Problem Statement, Our approach, Survey Goal, and Survey Questions. Problem Statement introduces the research background regarding what is architecture erosion and its violation symptoms, as well as why it is important to investigate violation symptoms. Our approach describes the approach that we developed to automatically identify violation symptoms with high performance from a large number of code reviews as well as two examples of violation symptoms of architecture erosion detected by our approach from code review comments. In Survey Goal, we explain the purpose of this survey. In Survey Questions, we ask five general questions and present six statements concerning the usefulness of our approach. Note that, before the formal survey, we did a pilot survey (receiving 6 responses) with the purpose of revising and improving our survey questions and examples.
@@
The five general survey questions include demographic information (e.g., their countries and work experience). The six statements are assessed using a five-point Likert scale (Strong Agree, Agree, Neutral, Disagree, and Strongly Disagree) with one option (“I don’t know”). This scale enables participants to rate the extent that they agree or disagree with the statements. Finally, at the end of the survey, we ask an open-ended question to let the respondents freely provide comments and suggestions about our survey. To formulate the statements, we took inspiration from the survey questions of two survey studies (Nasab2021ais, ; Khalajzadeh2022sda, ) that evaluate the usefulness of automatic classifiers based on ML and DL techniques in the software engineering community. We chose to use declarative statements in our validation survey since declarative statements are deemed appropriate for exploratory studies to convey simple and clear judgements; in contrast, descriptive statements are more suitable in explanatory studies to provide the details of the judgements in various aspects. The statements are as follows:
@@
Data Analysis. For the survey questions, we applied descriptive statistics to analyze the responses to the demographic and Likert scale questions. For the open-ended interview questions, interviews were recorded and transcribed; subsequently, we used the open coding and constant comparison method (Adolph2011CC, ) to analyze and categorize the qualitative responses.
4. Results
4.1. RQ1: Identifying Violation Symptoms